59,231.41J of heat is added to 1000 g of ice at -10 oC. How much ice melts?
Answer #1

First the ice must wam to 0

Q = mc(delta T)

Q = (1)(2090)(10) = 20900

The heat left = 59231.41 - 20900 = 38331.41

Then Q = mH

38331.41 = m(3.3 X 105)

m = .116 kg (116 grams)

(This is done using the values from my book. If this is slightly off, comment below with values used by your book and a correct ion can be made)